# Diffusion Models for Reinforcement Learning: A Survey

### 11 Dec 2023 v2

@@ 3.3 Guided Sampling Methods
Diffusion models with guided sampling methods care about the conditioned data distribution $p(x|y)$, which makes it possible to generate samples with attributes of the label $y$. Based on whether an extra classifier model is to be trained, the methods are divided into two categories: classifier guidance and classifier-free guidance. An advantage of classifier guidance sampling is that the classifier and the diffusion model are trained independently. If you already have a diffusion model, you just need to train a classifier and integrate it with your diffusion model while sampling. Classifier-free guidance retrains the model totally, but it is more expressive and performs better compared to classifier guidance sampling methods.
3.3.1 Classifier Guidance
A simple thought is to train a differentiable discriminative model that presents $p(y|x)$. In other words, all we need to do is to train an extra classifier $p(y|x^{t})$ trained on noisy samples $x^{t}$. Assume the classifier has been pre-trained. Then, the reversing process is described as
$p_{\theta,\phi}(x^{t}|x^{t+1},y)=Zp_{\theta}(x^{t}|x^{t+1})p_{\phi}(y|x^{t})~{},$ (8)
where $Z$ is a normalization factor. Dhariwal and Nichol (2021) state that Eq. (8) can be approximately regarded as another Gaussian distribution:
$p(x^{t}|x^{t+1},y)=\mathcal{N}(\mu(x^{t},t)+s\Sigma(x^{t},t)g,\Sigma(x^{t},t))~{},$
where $g=\nabla_{x^{t}}\log p_{\phi}(y|x^{t})|_{x^{t}=\mu}$ and $s$ is the guidance scale to control the effect of conditions. The $\mu(x^{t},t)$ and $\Sigma(x^{t},t)$ function is identical to those in DDPM (Eq. (4) and Eq. (3)). Therefore, the sampling process can be explicitly presented as follows:
$x^{t+1}\sim\mathcal{N}(\mu+s\Sigma g,\Sigma)~{},$
where $\Sigma$ is short for $\Sigma(t,x^{t})$.
3.3.2 Classifier-free Guidance
Rather than estimating the conditioned data distribution $p(x|y)$ directly, classifier-free methods try to predict the score function $\nabla_{x}\log p(x|y)$. With the Bayes Theorem, we can decompose the score function into an unconditional term and a classifier conditioning term as
$\nabla_{x}\log p(x|y)=\nabla_{x}\log p(y|x)+\nabla_{x}\log p(x)~{}.$ (9)
Unlike the classifier guidance, the original training setup is modified so the diffusion model should be retrained. The network to estimate noise is defined as $\epsilon(x^{t},c)$ ($c=\varnothing$ for unconditional inputs). In practice, the conditional and unconditional models are trained with the same set of network parameters by randomly setting $c=\varnothing$ with a pre-specified probability during training. From Eq. (9) we can infer
$\nabla_{x}\log p(y|x)=\nabla_{x}\log p(x|y)-\nabla_{x}\log p(x)~{}.$
As Song et al. (2022) reveal, diffusion models and the score functions are equivalent, which indicates $\nabla_{x^{t}}\log p(x^{t})\propto\epsilon(x^{t},t)$. As a result, we can now substitute $\nabla_{x}\log p(y|x)$ with this into the formula for classifier guidance
$\bar{\epsilon}_{w}(x^{t},y)$ $=\epsilon_{\theta}(x^{t},y)+w(\epsilon_{\theta}(x^{t},y)-\epsilon_{\theta}(x^{t}))$
$=(1+w)\epsilon_{\theta}(x^{t},y)-w\epsilon_{\theta}(x^{t})~{},$
where $w$ stands for the guidance scale.
@@ 3.4 Fast Sampling Methods
Diffusion model has been criticized for its prolonged iterative sampling time. Several fast sampling skills are proposed to solve this issue. Generally, the skills extend the diffusion model to a more fundamental paradigm to acquire efficiency. We categorize the methods into two categories: those that do not involve learning (learning-free) and those that require extra learning sessions (learning-based).
Learning-free sampling methods.
@@
Denoising Diffusion Implicit Models (DDIM) Song et al. (2022) is one of the earliest works on sampling acceleration. It aims at extending DDPM to a non-Markovian case by learning another Markov chain $q_{\theta}(x^{t-1}|x^{t},x^{0})$. The work also reveals that DDPM and DDIM are special cases of a more general paradigm. Moreover, it is observed in later works that DDIM is the discrete version of solving the PF-ODE Eq. (7). Then, some high-order solvers emerged, such as DPM-solver Lu et al. (2022), which provides an excellent trade-off between sample quality and sampling speed. With DDIM as its first-order version, DPM-solver boosts the efficiency of solving PF-ODE, outperforming common numerical ODE solvers like Runge-Kutta. Consequently, DPM-solver has become one of the most frequently used fast sampling methods.
Learning-based sampling methods.
@@
Learning-based sampling is another approach to fast sampling. Unlike the learning-free methods, they include extra learning processes to reach a higher sampling efficiency at a slight expense of sampling quality. A recent work, Truncate Diffusion Probabilistic Model (TDPM) Zheng et al. (2023), proposes that both the diffusion and denoising process can be truncated so that the iterative steps are reduced. Specifically, the forward process is truncated when the sample is noisy enough, and the denoising process starts from a relatively noisy sample (not pure noise), which will be learned by a network or else. Moreover, Watson et al. (2021) learn a strategy to select the best $K$ time steps to maximize the training objective for the DDPMs, which also decreases the denoising steps.
@@
AdaptDiffuser Liang et al. (2023) Offline Classifier
@@
Crossway Diffusion Li et al. (2023b) Offline; Robotics None
SGP Suh et al. (2023) Offline; Robotics None
GSC Mishra et al. (2023) Imitation; Robotics Classifier
UniPi Du et al. (2023a) Imitation; Robotics Classifier-free
ChainedDiffuser Xian et al. (2023) Imitation; Robotics None
@@
MTDiff-p He et al. (2023a) Offline; Multi-task Classifier-free
MetaDiffuser Ni et al. (2023) Offline; Multi-task Classifier-free
SafeDiffuser Xiao et al. (2023) Offline; Safe Self-proposed
MADiff Zhu et al. (2023) Offline; Multi-agent Classifier-free
HDMI Li et al. (2023a) Offline; Hierarchical Classifier-free
@@
UniSim Yang et al. (2023) Trajectory Generation Classifier-free
@@
SfBC Chen et al. (2023a) Offline Policy Gradient
IDQL Hansen-Estruch et al. (2023) Offline Policy Gradient
EDP Kang et al. (2023) Offline Policy Gradient
DiffCPS He et al. (2023b) Offline None
NoMaD Sridhar et al. (2023) Robotics None
BESO Reuss et al. (2023) Offline; Goal-conditioned Classifier-free
CEP Lu et al. (2023a) Offline; Image Synthesis Policy Gradient
@@
CPQL Chen et al. (2023d) Offline; Online Q-loss
@@
MTDiff-s He et al. (2023a) Offline; Multi-task Classifier-free
@@
LatentDiffuser Li (2023) Offline Policy Gradient
 Hegde et al. (2023) Quality Diversity None
@@
4 The Roles of Diffusion Models in RL
Diffusion models have proven their ability to generate diverse data and model multi-modal distributions. Considering the long-existing challenges introduced in Section 2, it is sufficient to improve the performance and sample efficiency of RL algorithms with diffusion models. In Fig. 1, we illustrate how diffusion models play a different role in RL compared to previous solutions. Current works applying diffusion models on RL mainly fall into three categories: using diffusion models as the planner, as the policy, and as the data synthesizer. It is essential to note that we include methods that generate action-only sequences as planners, even though some of the representative works have “policy” in their names, e.g., Diffusion Policy Chi et al. (2023). Generating multi-step action sequences can be viewed as planning in action space, and the use of diffusion models to ensure temporal consistency is similar to other planning-based diffusion methods. The following subsections will illustrate overall frameworks and representative papers for each category.
4.1 Planner
Planning in RL refers to the process of using a model of the environment to make decisions imaginarily, and then selecting the best action in order to maximize a cumulative reward signal. This process usually simulates or explores different sequences of actions and states, predicting the outcomes of its decisions, thus resulting in better actions from the perspective of a longer horizon. Therefore, planning is commonly applied in the MBRL framework. However, the decision sequences used for planning are generated autoregressively, which may lead to severe compounding errors, especially in the offline setting, due to the limited data support. Diffusion models provide a possible solution since they can generate the whole sequence simultaneously.
A general framework of diffusion models as planners is shown in Fig. 2(a), which was first proposed by Janner et al. (2022). Inputs and outputs of the diffusion model are usually clips of the real trajectory $\tau=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$, denoted as $x(\tau)=(e_{1},e_{2},\ldots,e_{H})$. Here $e_{t}$ represents for the selected elements from $(s_{t},a_{t},r_{t})$, where various choices can be made as $e_{t}=(s_{t},a_{t})$ Janner et al. (2022); Liang et al. (2023); He et al. (2023a); Xiao et al. (2023), or $e_{t}=(s_{t},a_{t},r_{t})$ He et al. (2023a); Hu et al. (2023), $e_{t}=s_{t}$ Ajay et al. (2023); Zhu et al. (2023), or $e_{t}=a_{t}$ Chi et al. (2023); Li et al. (2023b). $y(\tau)$ is the guidance that contains the desired property of the generated trajectory, such as the discounted return, the indicator of whether the task is completed, etc. Such guidance is meant to lead the diffusion model to generate good trajectories instead of those that only satisfy the environment dynamics. Suppose the diffusion model $\epsilon_{\theta}$ parameterized by $\theta$ predicts the noise added in the forward process, and $x^{k}\coloneqq x(\tau^{k})$ is the corresponding clip of $\tau$ at the $k$-th diffusion step. Either classifier guidance or classifier-free guidance is incorporated. For classifier guidance Janner et al. (2022); Liang et al. (2023), $\epsilon_{\theta}$ is trained as usual by taking $\tau^{k}$ and the timestep $k$ as input:
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon}[\|\epsilon-\epsilon_{\theta}(x^{k},k)\|_{2}^{2}]~{}.$
After training, suppose $\mu_{\theta}(x^{k},k)$ and $\Sigma$ are the mean and the variance of the original reverse process $p_{\theta}(x^{k-1}|x^{k})$, respectively. The gradient of $y(\tau)$ is injected during sampling as the guidance:
$p_{\theta}(x^{k-1}|x^{k},y(\tau))\approx\mathcal{N}(x^{k-1};\mu_{\theta}+\alpha\Sigma^{k}\nabla y(\tau),\Sigma^{k})~{},$
where $\alpha$ is a hyperparameter. In contrast, classifier-free guidance Ajay et al. (2023); He et al. (2023a); Hu et al. (2023) injects $y(\tau)$ in both training and sampling stages. The noise prediction model $\epsilon_{\theta}$ now takes an additional $y(\tau)$ as input, giving the training objective:
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,k,\epsilon,\beta}[\|\epsilon-\epsilon_{\theta}(x^{k},(1-\beta)y(\tau)+\beta\varnothing,k)\|_{2}^{2}]~{},$
where $\beta$ is sampled from a Bernoulli distribution with probability $p$. At sampling, $y(\tau)$ is directly fed into $\epsilon_{\theta}$, as
$p_{\theta}(x^{k-1}|x^{k},y(\tau))\approx\mathcal{N}(x^{k-1};\mu_{\theta}(x^{k},y(\tau),k),\Sigma^{k})~{}.$
As the time consumption of sampling using the original DDPM is unacceptable when deploying RL in the online environment, fast sampling methods are commonly incorporated, and the diffusion step for sampling is usually set to 100 or 200. Besides, to guarantee sample stability during the sampling process, the first several steps of a sampled trajectory can be replaced with the ground truth as a hard constraint. For example, when planning the trajectory after $e_{1:h}$, the model first generates the full trajectory $\tilde{x}^{k}=\tilde{e}_{1:H}^{k}$ at the $k$-the diffusion step. Then the first $h$ steps are replaced, forming $\tilde{x}^{k}\leftarrow(e_{1},e_{2},\ldots,e_{h},\tilde{e}_{h+1}^{k},\ldots,\tilde{e}_{H}^{k})$.
The core ideas behind why non-autoregressive diffusion models can generate Markovian decision sequences are stated by Janner et al. (2022). With the U-Net backbone, the sampled token $e_{t}^{k}$ takes information in a local receptive field $e_{t-l:t+l}^{k+1}$, where $l$ is determined by the model structure. Once the local consistency is guaranteed, the global consistency will also be built during the denoising process as information spreads among the trajectory. Though this raises new confusion that using future information is anti-causal for sequential problems, RL includes a hidden hypothesis that the current decision will lead to optimal future outcomes. That is, utilizing future information does not violate the general RL framework.
Benefiting from the non-autoregressive scheme, diffusion models surpass traditional models in generating planning sequences. First, such generations do not suffer from the aforementioned compounding errors. Then, diffusion models fit for both continuous-reward and sparse-reward settings Janner et al. (2022), even for more challenging multitask planning He et al. (2023a), safe planning Xiao et al. (2023), and generating based on multi-modal observations Du et al. (2023a); Chi et al. (2023). Besides, diffusion models can also learn from non-Markovian data by exploiting temporal information Hu et al. (2023).
4.2 Policy
Compared with traditional RL taxonomy, which roughly divides RL algorithms into MBRL and model-free RL, using diffusion models as the planner is similar to MBRL and focuses on capturing the environment dynamics. In contrast, taking diffusion models as the policy follows the framework of model-free RL. Section 2.1 states the main drawbacks of offline policy learning frameworks: over-conservatism and poor capability on diversified datasets. With excellent expressive ability on multi-modal distribution, many works utilize diffusion models as the policy to tackle these problems.
Diffusion-QL Wang et al. (2023) first explores the advantages of the diffusion policy in offline RL and finds that it can perfectly fit on dataset generated by strong multi-modal behavior policies, where previous distance-based BC methods fail. Experiment results also support the dominance of diffusion policies on diversified datasets Lu et al. (2023a); Chen et al. (2023a); Hansen-Estruch et al. (2023); Ada et al. (2023); Li et al. (2023c). Compared with diffusion models as planners, the diffusion target of the diffusion policy is simply the action given the current state, as shown in Fig. 2(b). Suppose the noise predictor is $\epsilon_{\theta}(a^{k},k,s)$ parameterized by $\theta$, and the derived diffusion model is $\mu_{\theta}(a|s)$. To guide the model sampling actions with high Q-values, it is necessary to take $Q(s,a)$ into consideration. Some papers Chen et al. (2023a); Lu et al. (2023a); Hansen-Estruch et al. (2023); Kang et al. (2023) construct the policy by (advantage) weighted regression as
$\pi_{\theta}(a|s)\propto\mu_{\theta}(a|s)\exp(\alpha Q(s,a))~{},$
where $\alpha$ is a hyperparameter. Following this, Chen et al. (2023a) decouple the policy learning into behavior learning and action evaluation, which allows more freedom in the choice of guidance. It also proposes in-sample planning for Q-learning, which avoids extrapolation errors in previous offline RL methods. Lu et al. (2023a) further generalize this framework to any energy reweighted distribution $p(x)\propto q(x)\exp(-\beta\mathcal{E}(x))$, where $\mathcal{E}(x)$ is an energy function, and train it via contrastive learning. Approaches using explicit Q-values as guidance are also considered. Wang et al. (2023) and Ada et al. (2023) subtract a weighted expectation of $Q(s,a)$ in the training loss as
where $\eta$ is a hyperparameter, and $\mathcal{D}$ is the offline dataset.
Fast reaction is crucial when deploying policies in online environments. Therefore, almost all diffusion policies use smaller diffusion steps during sampling, usually about 15 steps. ODE solvers such as the DPM-solver Lu et al. (2022) are also used to accelerate sampling Chen et al. (2023a); Lu et al. (2023a); Kang et al. (2023); Li et al. (2023c)Pearce et al. (2023) compare three different backbone structures of diffusion models, finding that MLP-sieve or Transformer is also sufficient in expressiveness. Kang et al. (2023) introduce action approximation, which allows one-step action sampling in the training stage.
4.3 Data Synthesizer
In addition to fitting multi-modal distributions, a simple and common use of diffusion models is to generate more training samples, which has been widely applied and proven in computer vision. Therefore, it is natural to apply the diffusion model as a data synthesizer on RL datasets, as data scarcity is a practical challenge of RL, as stated in Section 2.2. To guarantee consistency of synthetic data to the environment dynamics, previous data augmentation approaches in RL usually add small perturbations to existing states and actions Sinha et al. (2021). In contrast, Fig. 2(c) illustrates that diffusion models learn the data distribution from the whole dataset $\mathcal{D}$, and enable generating highly diversified data while keeping consistency. Lu et al. (2023b) investigate the ability of diffusion models as the data synthesizer in both offline and online settings. It directly trains the diffusion model from the offline dataset or the online replay buffer and then generates more samples for policy improvement. Analysis shows that the quality of data generated by diffusion models is higher than those generated by explicit data augmentation in diversity and accuracy. With synthetic data, the performance of the offline policy and the sample efficiency of the online policy are significantly improved. He et al. (2023a) deploy diffusion models to augment data for multitask offline datasets and achieves better performance than those on single-task datasets. It claims that fitting on multiple tasks may enable implicit knowledge sharing across tasks, which also benefits from the multi-modal property of diffusion models.
4.4 Others
Besides directions that have been mainly focused on, some ways of improving RL with diffusion models are under exploration. Hegde et al. (2023) take a similar idea as hyper networks in meta-learning, generating parameters of policies for quality diversity RL. The trained diffusion models compress the parameters of various policies into the latent space while maintaining the ability to generate policies under certain conditions. Mazoure et al. (2023) estimate value functions with diffusion models by learning the discounted state occupancy, combined with a learned reward estimator. Then, the value function can be directly computed by definition, where future states are sampled from the diffusion model. Venkatraman et al. (2023) follow Latent Diffusion Models Rombach et al. (2022) by first encoding the high-level trajectories into semantically rich representations, then applying diffusion models on them. Conditioning on latent representations, Q-functions, and policies achieves higher capability without significant extrapolation errors.

### 23 Feb 2024 v4

@@ 3.3 Guided Sampling Methods
Guided sampling from diffusion models considers sampling from the conditional distribution $p(x|y)$, where $y$ is the desired attribute of generated samples. Two main categories of guidance are classifier-guidance and classifier-free guidance.
Classifier guidance.
Classifier guided sampling relies on a differentiable classifier model $p_{\phi}(y|x)$. Specifically, since the guidance needs to be performed on each denoising step, the classifier model $p(y|x^{t})$ is trained on noisy samples of $x$ and corresponding attribute $y$. The conditional reverse process can be written as
$p_{\theta,\phi}(x^{t-1}|x^{t},y)=Zp_{\theta}(x^{t-1}|x^{t})p_{\phi}(y|x^{t-1})~{},$ (4)
where $Z$ is the normalization factor. Dhariwal and Nichol (2021) approximate Eq. (4) by another Gaussian distribution:
$p_{\theta,\phi}(x^{t-1}|x^{t},y)=\mathcal{N}(\mu^{t}+w\Sigma^{t}g,\Sigma^{t})~{},$ (5)
where $g=\nabla_{x^{t}}\log p_{\phi}(y|x^{t})|_{x^{t}=\mu^{t}}$ and $w$ is the guidance scale to control the strength of conditions. $\mu^{t}$ and $\Sigma^{t}$ are the mean and the covariance matrix in Eq. (2), respectively.
Classifier-free guidance.
Classifier-free sampling relies on an extra conditional noise model $\epsilon_{\theta}(x^{t},y,t)$. In practice, the conditional and unconditional models share the same set of parameters, and the unconditional model is represented by setting $y$ as a dummy value $\varnothing$. Ho and Salimans (2022) state that the noise learning target in Eq. (3) is a scaled score function of $p(x^{t})$, i.e., $\epsilon(x^{t},t)=-{\sigma^{t}}\nabla_{x^{t}}\log p(x^{t})$ and $\sigma^{t}=\sqrt{\beta_{t}}$. By using Bayes theorem, we have
$\nabla_{x^{t}}\log p(y|x^{t})=-1/\sigma^{t}(\epsilon(x^{t},y,t)-\epsilon(x^{t},t))~{}.$
According to Eq. (5), we can derive the guided noise predictor as $\bar{\epsilon}_{\theta}(x^{t},y,t)=\epsilon_{\theta}(x^{t},t)-w\sigma^{t}\nabla_{x^{t}}\log p(y|x^{t})$. Replacing the score function with the noise model predictions, the noise used in classifier-guided sampling can be written as
$\hat{\epsilon}_{w}(x^{t},y,t)=w\epsilon_{\theta}(x^{t},y,t)+(1-w)\epsilon_{\theta}(x^{t},t)~{}.$
@@ 3.4 Fast Sampling Methods
Various fast sampling methods are proposed to overcome the prolonged iterative sampling time of diffusion models. We summarize these methods into two categories: those that do not involve learning and those that require extra learning, and describe representative works in each category.
Learning-free methods.
@@
DDIM Song et al. (2022) is one of the seminal works on sampling acceleration. It extends DDPM to a non-Markovian formulation by learning another Markov chain $q_{\theta}(x^{t-1}|x^{t},x^{0})$. Some high-order solvers are proposed for diffusion sampling, such as DPM-solver Lu et al. (2022), which provides an excellent trade-off between sample quality and sampling speed. With DDIM as its first-order version, DPM-solver boosts the efficiency of solving PF-ODE, outperforming common numerical ODE solvers.
Learning-based methods.
@@
Learning-based sampling methods require extra training to obtain a higher sampling efficiency at a slight expense of sampling quality. A recent work, Truncate Diffusion Probabilistic Model (TDPM) Zheng et al. (2023), demonstrates that both the noising and denoising process can be early terminated to reduce the iterative steps. Moreover, Watson et al. (2021) propose a strategy to select the best $K$ time steps to maximize the training objective for the DDPMs, which also decreases the denoising steps.
4 The Roles of Diffusion Models in RL
Fig. 1 illustrates how diffusion models play a different role in RL compared to previous solutions. Current works applying diffusion models on RL mainly fall into three categories: as planners, as policies, and as data synthesizers. It is essential to note that we include methods that generate action-only sequences as planners, even though some of the representative works have “policy” in their names, e.g., Diffusion Policy Chi et al. (2023). Generating multi-step action sequences can be viewed as planning in action space, and the use of diffusion models to ensure temporal consistency is similar to other planning-based diffusion methods. The following subsections will illustrate overall frameworks and representative papers for each category.
4.1 Planner
Planning in RL refers to using a dynamic model to make decisions imaginarily and selecting the appropriate action to maximize cumulative rewards. This process usually explores various sequences of actions and states, thus improving decisions over a longer horizon. Planning is commonly used in the MBRL framework with a learned dynamic model. However, the planning sequences are usually simulated autoregressively, which may lead to severe compounding errors, especially in the offline setting due to limited data support. Diffusion models offer a promising alternative as they can generate multi-step planning sequences simultaneously.
A general framework of diffusion planners is shown in Fig. 2(a). Diffusion models are designed to generate clips of the trajectory $\tau=(s_{1},a_{1},r_{1},\ldots,s_{H},a_{H},r_{H})$, denoted as $x(\tau)=(e_{1},e_{2},\ldots,e_{H})$. $H$ is the planning horizon. Here $e_{t}$ represents the selected elements from $(s_{t},a_{t},r_{t})$, where various choices can be made as $e_{t}=(s_{t},a_{t})$ Janner et al. (2022); Liang et al. (2023a); He et al. (2023a), $e_{t}=(s_{t},a_{t},r_{t})$ He et al. (2023a); Hu et al. (2023), $e_{t}=s_{t}$ Ajay et al. (2023); Zhu et al. (2023), or $e_{t}=a_{t}$ Chi et al. (2023); Li et al. (2023b). RL datasets often contain trajectory data of varying quality. In order to make the diffusion planner generate high rewarded trajectories during evaluations, guided sampling techniques are widely adopted. The guidance can be either injected in the sampling stage following the classifier guided sampling or in both the training and sampling stages following the classifier-free sampling.
When deploying the trained diffusion planner for online evaluations, fast sampling methods are usually adopted to reduce the inference time. Besides, to ensure the planned trajectory is congruous with the agent’s current state, before each denoising step, the first $h\,(h\geq 1)$ steps of the noisy trajectory are substituted with the $h$ steps of historical states observed by the agent. Here, $h$ is a hyperparameter where a larger $h$ can better handle partially observable and non-Markovian settings but increases the modeling complexity.
4.2 Policy
Based on whether they rely on a dynamic model for making decisions, RL algorithms can be categorized into MBRL and model-free RL. Under such classification criteria, using diffusion models as planners is more akin to MBRL, as the generated trajectories encapsulate dynamics information. Another perspective is that diffusion planner can be seen as a combination of policy and dynamic model Janner et al. (2022). In contrast, using diffusion models as policies focuses on improving existing model-free RL solutions. Section 2.1 states the main drawbacks of current offline policy learning methods: over-conservatism and lack of expressiveness. Many works use diffusion models as the policy class in model-free RL to tackle these problems.
Diffusion-QL Wang et al. (2023) first combines the diffusion policy with the Q learning framework and finds that it can perfectly fit on datasets collected by strong multi-modal behavior policies, where previous distance-based policy regularization approaches fail. Compared with using diffusion models as planners, the diffusion target of the diffusion policy is simply the action given the current state, as shown in Fig. 2(b). Suppose the noise predictor is $\epsilon_{\theta}(a^{k},k,s)$ parameterized by $\theta$, and the derived action mean is $\mu_{\theta}(a|s)$. To guide the model sampling actions that can lead to high returns, it is necessary to take $Q(s,a)$ into consideration. Diffusion-QL includes a weighted Q maximization term into the diffusion training loss as
where $\eta$ is a hyperparameter, and $\mathcal{D}$ is the offline dataset. Some works Chen et al. (2023a); Lu et al. (2023a); Hansen-Estruch et al. (2023); Kang et al. (2023) construct the policy by (advantage) weighted regression as
$\pi_{\theta}(a|s)\propto\mu_{\theta}(a|s)\exp(\alpha Q(s,a))~{},$
where $\alpha$ is the temperature hyperparameter. Following this, Chen et al. (2023a) decouple the policy learning into behavior learning and action evaluation, which allows more freedom in the choice of guidance. They also propose in-sample planning for Q-learning, avoiding extrapolation errors in previous offline RL methods. CEP Lu et al. (2023a) further extends this framework to sample from the more general energy-guided distribution $p(x)\propto q(x)\exp(-\beta\mathcal{E}(x))$. Here, $\mathcal{E}(x)$ is an energy function, and in the RL setting, it is trained via contrastive learning to match the in-sample softmax of Q functions. Since there are off-the-shelf Q functions or energy functions after training, some methods use those functions to further improve the sampling quality during evaluations. They first sample multiple candidate actions for a given state and use Q or energy values to perform weighted sampling or just select the best candidate.
Fast reaction is crucial when deploying policies in online environments. Therefore, almost all diffusion policies use smaller diffusion steps during sampling, usually about 15 steps. ODE solvers such as the DPM-solver Lu et al. (2022) are also used to accelerate sampling Chen et al. (2023a); Lu et al. (2023a); Kang et al. (2023); Li et al. (2023c). Kang et al. (2023) introduce action approximation, which allows one-step action sampling in the training stage.
4.3 Data Synthesizer
In addition to fitting multi-modal distributions, a simple and common use of diffusion models is to create synthetic data, which has been widely applied in computer vision. Therefore, the diffusion model is a natural data synthesizer for RL datasets because data scarcity is a practical concern. To ensure consistency of synthetic data to the environment dynamics, previous data augmentation approaches in RL usually add minor perturbations to states and actions Sinha et al. (2021). In contrast, Fig. 2(c) illustrates that diffusion models generate diverse and consistent data by learning the data distribution from the entire dataset $\mathcal{D}_{\mathrm{real}}$. The diffusion model first learns the parameterized data distribution $\rho_{\theta}(\tau)$ from the real data $\mathcal{D}_{\mathrm{real}}$, and generates desired synthetic data by
$\mathcal{D}_{\mathrm{syn}}=\{\tau\sim\rho_{\theta}(\tau)\}~{}.$
Then, real and synthetic data are combined together as
$\mathcal{D}=\mathcal{D}_{\mathrm{real}}\cup\mathcal{D}_{\mathrm{syn}}~{},$
and $\mathcal{D}$ is used for policy learning. In online settings, the policy interacts with the environment, collects more real data into $\mathcal{D}_{\mathrm{real}}$, and updates the diffusion model. As a result, the diffusion model and the policy are updated alternately.
4.4 Others
Besides the primary directions discussed above, other ways of improving RL with diffusion models are less explored. Mazoure et al. (2023) estimate value functions with diffusion models by learning the discounted state occupancy, combined with a learned reward estimator. Then, the value function can be directly computed by definition, where future states are sampled from the diffusion model. Venkatraman et al. (2023) first encode the high-level trajectories into semantically rich latent representations, then apply diffusion models to learn the latent distribution. Conditioning on latent representations improves the capability of Q-functions and policies without significant extrapolation errors. Rigter et al. (2023) use a diffusion dynamic model and allow an online RL policy to collect synthetic trajectories on it. The interplay of the diffusion dynamic model and RL policy is done by alternating between state denoising and Langevin dynamics of policy actions.
@@ 4.4 Others
AdaptDiffuser Liang et al. (2023a) Offline Classifier
@@
LatentDiffuser Li (2023) Offline Energy Function
HDMI Li et al. (2023a) Offline; Hierarchical Classifier-free
SafeDiffuser Xiao et al. (2023) Offline; Safe None
MADiff Zhu et al. (2023) Offline; Multi-agent Classifier-free
MTDiff-p He et al. (2023a) Offline; Multitask Classifier-free
MetaDiffuser Ni et al. (2023) Offline; Multitask Classifier-free
@@
SkillDiffuser Liang et al. (2023b) Imitation; Multitask; Hierarchical Classifier-free
@@
UniSim Yang et al. (2023b) Trajectory Generation Classifier-free
@@
EDP Kang et al. (2023) Offline Q-loss
SfBC Chen et al. (2023a) Offline Sample & Reweight
IDQL Hansen-Estruch et al. (2023) Offline Sample & Reweight
DiffCPS He et al. (2023b) Offline Convex Optimization
CPQL Chen et al. (2023d) Offline; Online Q-loss
CEP Lu et al. (2023a) Offline; Image Synthesis Energy Function
@@
NoMaD Sridhar et al. (2023) Imitation; Robotics None
BESO Reuss et al. (2023) Imitation; Goal-conditioned Classifier-free
@@
MTDiff-s He et al. (2023a) Offline; Multitask Classifier-free
@@
PolyGRAD Rigter et al. (2023) Dynamic Model Online None